A student weighs out a 2.546 g sample of solid CoCO3●nH2O. Upon heating the following reaction occurs:

CoCO3●nH2O(s)CoO(s) + CO2(g) + n H2O(g) After the reaction has gone to completion the mass of solid

remaining is 0.840 g.

a. How many moles of CO2 gas were formed? You must show your work to receive any credit.

b. What is the molecular weight of CoCO3●nH2O?

c. What is the value of “n”? Round to the nearest whole number. You must show your work to receive any credit.
Answer #1

a) Initial mass of solid is reactant = 2.546 g

Final mass of solid after reaction is CoO = 0.84 g

Molar mass of CoO = 75 g/mol

Moles of CoO = mass/molar mass = 0.84/75 = 0.0112 moles.

According to reaction stoichiometry, when one mole of CoO is formed, one mole of CO2 is formed.

When 0.0112 moles of CoO is formed, moles of CO2 formed = 0.0112 moles

Thus, moles of CO2 formed = 0.0112 moles. Mass of CO2 = molar mass of CO2 x moles of CO2 = 44 x 0.0112 = 0.4928 g

b) Using mass conservation, mass of reactant = mass of product = 2.546 g.

for one mole of CoO to form, one mole of CoCO3.nH2O is required. Therefore, for 0.0112 moles of CoO to form, moles of CoO3 required = 0.0112.

Molar mass of CoO3.nH2O = mass of CoO3.nH2O/moles of CoO3.nH2O = 2.546/0.0112 = 227.32 g/mol

Thus, molar mass/molecular weight of CoCO3.nH2O = 227.32 g/mol

c) Atomic weight of Co = 59 a.m.u, Atomic weight of C = 12 a.m.u and Atomic weight of O = 16 a.m.u

Therefore, molar mass of CoCO3 = 59 + 12 + (3x16) = 119.

But molecular weight of CoCO3.nH2O = 227.32. The difference in weight is mass of nH2O = 227.32 - 119 = 108.32.g

value of n = (difference in weight of CoCO3.nH2O and CoCO3)/molar mass of H2O = 108.32/18 = 6.017 ~ 6

Thus, value of n = 6